Obtain all the basic solution of the system of linear equations
1
Expert's answer
2020-06-16T19:14:55-0400

As per the given question,

Let there are m linear equation which has n variables,

"\begin{cases}
 a_{11}x_1+a_{12}x_2.........+a_{1n}x_n=b_1 \\
 a_{21}x_1+a_{21}x_2.........+a_{2n}x_n=b_2 \\
a_{31}x_1+a_{32}x_3.........+a_{3n}x_n=b_3 \\.
..\\
...\\.
..\\
a_{n_1 1}+a_{n2}x_2.........+a_{nn}x_n=b_n \\
\end{cases}"


Let it is Ax=b

it have the basic feasible solution if,

"Ax=b ; x \geq0"

The Simplex Method uses the pivot procedure to move from one BFS to an “adjacent” BFSwith an equal or better objective function value.

Pivot Procedure:

  1. Choose a pivot element "a_{ij}"
  2. Divide row i of the augmented matrix "[A|b]"  by "a_{ij}"

"\begin{bmatrix}
...& a_{ij}&... & a_{il}&... \\
...& ...&... &...&...\\
 ...&...a_{kj}&...&a_{kl}&...\\
\end{bmatrix}\rightarrow \begin{bmatrix}
...& 1&... & \frac{a_{il}}{a_{ij}}&... \\
...& ...&... &...&...\\
 ...&...a_{kj}&...&a_{kl}&...\\
\end{bmatrix}"


3.For each row k(other than row i), add "\u2212a_{kj}x" row i to row k.

The element in row k, column l becomes "\u2212a_{kj}\u00d7a_il+a_kl"

"\begin{bmatrix}
...& 1&... & \frac{a_{il}}{a_{ij}}&... \\
...& ...&... &...&...\\
 ...&...a_{kj}&...&a_{kl}&...\\
\end{bmatrix} \rightarrow \begin{bmatrix}
...& 1&... & \frac{a_{il}}{a_{ij}}&... \\
...& ...&... &...&...\\
 ...&...0&...&a_{kl}-\frac{a_{kj}a_{il}}{a_{ij}}&...\\
\end{bmatrix}"